One carpenter can install 625 SF of ceiling tile in grid on one workday. How many hours of labor are needed to install 1 S.F.?

a. How many minutes does it take to install 1 S.F. of ceiling tile?
Answer #1

625 sf of ceiling tile in 1 workday i.e 24 hours

1 sf of ceiling tile in ?

1*24/625 hrs = 0.0384 hrs ~0.04 hrs

a. Minutes = 0.0384*60 = 2.304 minutes.